OpenStudy (anonymous):

A storage box has a volume of 4368 in cubed. the length is 24 in and the height is 1 in more than the width. Find the height and width of the box. Solve by factoring

OpenStudy (lgbasallote):

im assuming this is a rectungular prism.. volume = lwh volume = 4368 length = 24 height = 1 + w w = w substitute then "solve by factoring"

OpenStudy (lgbasallote):

do you understand? or must i go on?

OpenStudy (anonymous):

Please go on

OpenStudy (anonymous):

I'm trying to understand the process of getting there

OpenStudy (lgbasallote):

like i said...substitute 4368 = 24(1+w)(w) ---- ---------- 24 24 182 = (1+w)(w) you now have the "factor"

OpenStudy (lgbasallote):

but you need to go on...distribute w.. w + w^2 = 182 w+ w^2 - 182 = 0 factor this out and equate to zero...